### v1

# GEMeX-ThinkVG: Towards Thinking with Visual Grounding in Medical VQA via Reinforcement Learning

Abstract
Medical visual question answering aims to support clinical decision-making by enabling models to answer natural language questions based on medical images. While recent advances in multi-modal learning have significantly improved performance, current methods still suffer from limited answer reliability and poor interpretability, impairing the ability of clinicians and patients to understand and trust model-generated answers. To address this, this work first proposes a Thinking with Visual Grounding (ThinkVG) dataset wherein the answer generation is decomposed into intermediate reasoning steps that explicitly ground relevant visual regions of the medical image, thereby providing fine-grained explainability. Furthermore, we introduce a novel verifiable reward mechanism for reinforcement learning to guide post-training, improving the alignment between the model’s reasoning process and its final answer. Remarkably, our method achieves comparable performance using only one-eighth of the training data, demonstrating the efficiency and effectiveness of the proposal. The dataset is available at https://huggingface.co/datasets/BoKelvin/GEMeX-ThinkVG.
1 Introduction
Figure 1: Interpretability comparison: (Top) Traditional answer with grounding provided by GEMeX: separate textual and visual prompts; (Bottom) Our integrated, more detailed thinking process for answer generation, which explicitly grounds evidence in specific regions (shown in the colored texts) of the medical image, i.e., anatomical areas that support inferences, thereby enhancing the understanding of questions and answers.
Medical Visual Question Answering (VQA) has emerged as a promising paradigm for supporting clinical decision-making by enabling models to answer natural language questions based on medical images [16]. Recent advances in multi-modal learning have led to significant performance improvements across a range of Medical VQA benchmarks [19, 13, 8, 6], especially in the era of large vision language models [17, 34, 3, 36]. However, despite these gains, existing methods remain limited in their ability to provide interpretable and trustworthy responses—an essential requirement for real-world deployment in clinical settings. In particular, most current approaches generate answers without revealing the underlying reasoning process or the specific image regions that inform their decisions. This lack of transparency undermines user trust and hinders the integration of such systems into clinical workflows.
A recent work GEMeX [21] introduced a new dataset in which each VQA triplet is paired with a textual reason and a corresponding visual region as an explanation, bridging this gap to some extent. However, the textual and visual prompts are independent and lack coherence. Moreover, the textual reason merely serves as an explanation of the answer rather than providing guidance on how to solve the question based on the image. Therefore, based on GEMeX, we propose a new dataset GEMeX-ThinkVG that enhances interpretability in Medical VQA through introducing a Thinking with Visual Grounding mechanism, as shown in Figure 1. Different from the traditional chain-of-thought reasoning [35] and recent approaches [24, 15] that rely solely on reinforcement learning to incentivize the thinking process for solving multimodal tasks, our ThinkVG explicitly decomposes the answer generation process into a sequence of intermediate steps, along with grounding corresponding visual evidence within the medical image. By linking reasoning steps to spatially localized regions, we offer clinicians a fine-grained explanation of how the model arrives at its conclusions.
While leveraging the GEMeX-ThinkVG dataset for Medical VQA enables the modeling of explicit thinking, our empirical findings indicate that supervised fine-tuning (SFT) alone is insufficient to fully unleash the reasoning capabilities of the large vision language models (LVLMs). To address this, we adopt a reinforcement learning (RL) framework for post-training that builds upon the SFT-tuned model. It introduces a novel verifiable reward mechanism to incentivize alignment between the reasoning trajectory and relevant visual and textual cues, thereby enhancing both the accuracy of intermediate thinking steps and final answer generation. In contrast to existing methods that require large-scale datasets, our approach is data-efficient, achieving comparable performance using only one-eighth of the full training data.
In summary, our contributions are threefold:
  • We introduce GEMeX-ThinkVG, a Thinking with Visual Grounding dataset for Medical VQA, which provides step-by-step reasoning grounded in specific anatomical regions of medical images, offering better interpretability for understanding the questions and answers.
  • We perform supervised fine-tuning on the LVLM using our GEMeX-ThinkVG dataset, and further apply reinforcement learning with a novel verifiable reward mechanism to incentivize reasoning abilities, enabling the generation of more accurate thinking paths and final answers.
  • From comprehensive experiments, we demonstrate the effectiveness of the proposed dataset and reward mechanism that empower LVLM to achieve comparable performance with only one-eighth of the training data.
2 Related Work
2.1 Medical VQA Dataset
Recent years have witnessed significant progress in medical visual question answering (VQA) through the development of specialized datasets targeting diverse clinical challenges. VQA-RAD [16] established the foundation with 3,000 question-answer pairs focused on radiology images. SLAKE [20] expanded the scope with over 14,000 manually annotated QA pairs spanning CT, MRI, and X-ray modalities. OmniMedVQA [8] further broadened coverage across multiple body regions and imaging modalities to enhance model generalization capabilities. For specialized tasks, PathVQA [6] offers 32,000 QAs on histopathology images. MIMIC-Diff-VQA [7] emphasizes differential diagnosis between paired X-rays. GEMeX [21] enhances VQA with both textual and visual explanations, facilitating better understanding of the answers for patients. However, none of them incorporate a reasoning process for problem-solving, which hinders patients and junior doctors from fully understanding the questions and answers. To fill this gap, we propose a thinking with visual grounding dataset, GEMeX-ThinkVG, in this work.
2.2 Reinforcement Learning
Recently, the growing popularity of reasoning-focused large language models such as GPT-o1 [10] and DeepSeek-R1 [5] has brought increasing attention to the underlying mechanisms of reinforcement learning (RL). This interest has spurred extensive research and led to remarkable advancements in tasks such as mathematics [23, 32], code generation [12, 33], and hallucination detection [37, 30]. A notable contribution is the Group Relative Policy Optimization (GRPO) algorithm proposed in DeepSeekMath [28], which significantly simplifies the training process. Unlike traditional RL algorithms like PPO [12] that require a critic model to evaluate policy performance, GRPO compares groups of candidate responses directly, eliminating the need for an additional critic. GRPO’s simplicity has inspired further exploration into the reasoning capabilities of large vision-language models (LVLMs) for multimodal and visual tasks [9, 22, 29]. In the medical domain, GRPO has also been applied to incentivize the reasoning abilities of LVLMs for tasks such as out-of-distribution detection [24] and visual question answering [15]. Unlike these works that merely elicit textual descriptions of visual content, we propose a visually grounded reasoning paradigm that provides improved interpretability.
3 Construction of GEMeX-ThinkVG
Figure 2: An illustration of the proposed uncertainty-driven multi-agent framework for generating the GEMeX-ThinkVG dataset. The left (orange) section depicts the data generation process, while the right (blue) section shows the uncertainty-driven quality assurance pipeline, i.e., the “not sure” option prompts the agent to engage in deeper reflection and more deliberate decision-making.
3.1 Task Definition
Unlike conventional chain-of-thought approaches used in general tasks such as VQA [31], which typically generate only textual rationales before outputting the final answer, we propose a multimodal-level thinking process with visual grounding tailored for medical VQA. As illustrated in Figure 1, it decomposes the reasoning process into several steps with visual attention to specific image regions, aiming to deliver more detailed explanations and improved interpretability for patients and junior doctors. Formally, given a question $q$ for image $v$, we hope to provide a multi-modal thinking process $T$ involving visual grounding to attain answer $a$ via
$T=\left(\{r_{i}\}_{i=1}^{N_{r}},\{t_{j}\}_{j=1}^{N_{t}}\right),$ (1)
where $\{r_{i}\}_{i=1}^{N_{r}}$ is a set of image regions and $\{t_{j}\}_{j=1}^{N_{t}}$ is a set of relevant textual descriptions.
3.2 An Uncertainty-Driven Multi-Agent Framework
As illustrated in Figure 2, we construct the ThinkVG dataset based on GEMeX [21], the largest Medical VQA dataset that provides detailed visual and textual explanations for each VQA triplet. To ensure the high quality of ThinkVG, we introduce an uncertainty-driven multi-agent framework, centered around DeepSeek-R1 [5], which also coordinates multiple agents and leverages an uncertainty mechanism to enhance the accuracy of the generated data.
3.2.1 Data Generation
In Stage 1, we primarily use DeepSeek-R1 to generate the thinking process for each VQA triplet due to its outstanding reasoning capabilities. However, as DeepSeek-R1 is a text-based LLM and lacks the ability to directly perceive visual inputs, we utilize the grounded reports from GEMeX as an intermediary. In GEMeX’s reports, each sentence is aligned with a specific anatomical region, enabling DeepSeek-R1 to simulate a multimodal reasoning process required for solving the tasks. To optimize the prompt design, we randomly sample 50 questions from each question type (e.g., open-ended) in the training set—resulting in 200 questions in total—to evaluate performance and guide prompt iteration. The final optimized prompt is presented in Table 1.
messages = [“role”:“user”, “content”:
f“‘Suppose you are viewing a CXR that shows the following: “The hilar contours are normal [visual location: bilateral hilar structures ([116, 112, 227, 182])] …”. Given the question: “YOUR QUESTION”, provide a detailed thinking process (around 100 words) including specific visual location (e.g., (region [x1,y1,x2,y2])) about how to solve this question with answer “ANSWER”. You must assume that you are viewing the CXR image rather than reading the textual findings, thus do not output words like “observe the report” or “from report” or “report states” or “given findings” or “provided findings” or “described findings”.””
Table 1: Given a question-answer pair, our tailored prompt guides DeepSeek-R1 to generate ThinkVG data.
3.2.2 Uncertainty-driven Quality Assurance
Although manual verification and prompt modification are applied, the generated data still suffers from quality issues. To address this, we design a pipeline which contains three steps (Stage 2 to Stage 4) to ensure quality: (1) We first introduce an uncertainty-aware mechanism that enables DeepSeek-R1 to self-reflect and evaluate the reliability of its own outputs. Specifically, after generating the ThinkVG response, we prompt DeepSeek-R1 with the question: “Do you trust your thinking process? A. Yes B. No C. Not sure”. This self-reflection with uncertainty choice “Not sure” allows the model to carefully reassess its previous reasoning, which has been shown to improve output accuracy [2]. We retain only those ThinkVG responses for which the model answers “A. Yes.”; (2) To avoid bias caused by relying on a single LLM for both data generation and evaluation [25], we introduce additional agents to assess the quality of the generated data. As shown in Stage 3, we employ three additional agents—Qwen-2.5-72B [26], OpenBioLLM-70B [^1] https://huggingface.co/aaditya/Llama3-OpenBioLLM-70B, and GPT-4o [1]—as external reviewers. These agents are given grounded reports, questions, and answers to assess the accuracy of the thinking process. Similar to the self-reflection stage, we introduce an “uncertain” option to provide the agents with greater flexibility for consideration and decision-making. This process helps eliminate incorrect reasoning paths generated by DeepSeek-R1 based on its internal knowledge. Then, we adopt a voting mechanism: the ThinkVG is retained only if all three agents unanimously select “A. Correct” option; (3) Finally, we apply post-processing to eliminate ThinkVG outputs with significant errors, such as incorrect location information or instances that quote the report or findings directly, rather than generating content from an observational perspective.
3.3 Radiologist Validation
After auto-generation, a radiologist randomly samples 200 examples from each of the four question types in the dataset for manual inspection. Cross-validation reveals no clinically inconsistent reasoning chains, demonstrating the effectiveness of the proposed framework. In total, we generate 206,071 ThinkVG instances for 21,994 images, all drawn from the GEMeX training set. The detailed distribution is presented in Table 2.
Figure 3: An example from the SFT and RFT stages: after SFT, the model learns to think carefully and incorporates grounding to specific pathological regions for visual evidence. During the RFT stage, since both the answer and the grounded regions are correct, the two reward scores are each 1.0.
Open-ended Closed-ended Single-choice Multi-choice
62,712 37,340 53,905 52,114
Table 2: Distribution of generated ThinkVG data across the four question types in GEMeX.
4 Fine-tuning on GEMeX-ThinkVG: Learning to Think with Visual Grounding in Medical VQA
4.1 Supervised Fine-tuning
After generating the ThinkVG data, we first train the large vision language model (LVLM) using Supervised Fine-tuning (SFT) to equip it with task-aware thinking capabilities. To generate the outputs required by GEMeX—which include both the final answer and its corresponding visual location (omitting the reasons, as we provide a more detailed preceding thinking path)—the model must produce structured outputs in two distinct parts. First, it should output a thinking trace that incorporates explicit visual grounding, enclosed within the <think></think> tags. Second, it should deliver a concise final response that includes the predicted “answer” and the summarized “visual location”, wrapped within a designated XML-like structure: <response> <answer></answer> <location></location></response>. This format ensures both interpretability and alignment with the GEMeX evaluation requirements. An illustrative example is shown in Figure 3.
4.2 Reinforcement Fine-tuning
Although SFT can significantly improve thinking capabilities, we have observed in practice that: (1) due to the complexity of the task, the model sometimes generates incorrect thinking process after SFT, for example, incorrect analysis of images, which may be because SFT is better at helping LVLMs memorize rather than truly understand  [4]; (2) inaccurate region localization within the thinking path can propagate errors to the final answer, ultimately leading to degraded performance. Therefore, we adopt Reinforcement Fine-tuning (RFT), specifically through Group Relative Policy Optimization (GRPO) [28], as a post-training strategy to further activate and enhance the model’s thinking capabilities with a newly proposed reward mechanism.
4.2.1 Group Relative Policy Optimization
Given a question $q$ from question set $Q$, GRPO first samples a set of outputs $\{o_{i}\}_{i}^{G}$ from the policy model $\pi_{\theta_{old}}$ during each learning iteration. Then it optimizes the LVLM $\pi_{\theta}$ by maximizing the following objective:
$\quad\quad\mathcal{J}_{\text{GRPO}}(\theta)&=\mathbb{E}_{q\sim P(Q),\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\Bigg{[}\\ \frac{1}{G}\sum_{i=1}^{G}\min\Bigg{(}\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)}A_{i},&\text{clip}\left(\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},1-\varepsilon,1+\varepsilon\right)A_{i}\Bigg{)}\\ &-\beta D_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\Bigg{]}\end{aligned}$
where $A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\ldots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\ldots,r_{G}\})}$ is the relative advantage over group rewards $\{r_{i}\}_{i}^{G}$, $D_{\text{KL}}$ is the KL divergence between $\pi_{\theta}$ and $\pi_{\text{ref}}$ (the model after SFT stage) to prevent catastrophic forgetting caused by over optimization, $\epsilon$ and $\beta$ are the PPO clipping hyper-parameter [27] and the coefficient for the KL penalty, respectively.
messages = [“role”:“user”, “content”:
f“‘We would like to request your feedback on the performance of two AI assistants in response to the user question displayed above. For your reference, the visual content in the image is represented with caption describing the same image. Please rate the accuracy (most important), relevance of their responses, considering both answer and reason (if any). Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance. Please output both scores and your reaon in JSON format { assistant1: score, assistant2: score, reason:your_reason }.”””
Table 3: Our designed prompt for Semantics Reward.
4.2.2 Reward Functions
We propose two reward functions, tailored for the two output parts of GEMeX: answer and involved visual locations. (1) Semantics Reward: In GEMeX, four types of questions are defined (Table 2). For choice-based and closed-ended questions, reward design is straightforward, as we can directly use Accuracy Reward [24] to evaluate whether the model’s outputs are correct. Nevertheless, designing a reward function for open-ended questions is particularly challenging due to the absence of fixed answers. Unlike [18], which employs traditional machine learning metrics such as BLEU as rewards—metrics that have been shown to poorly capture essential medical content [11]—we introduce a semantics-aware accuracy reward. It enables the integration of all four types into a unified, coherent reward paradigm. Specifically, we input both the ground truth (GT) answer and the model-generated answer into OpenBioLLM-70B, which can provide two accuracy scores at the semantics level. When these two scores are close, i.e., if the absolute difference between them is less than 2, we assign a reward of 1 point; otherwise, the reward is 0. The specific prompt is shown in Table 3. (2) Grounding Reward: Besides giving answers, the model needs to provide the visual locations involved in solving the questions. Therefore, we first check whether the number of bounding boxes predicted by the model matches the number in the GT. If they match, we then evaluate whether the mean Intersection over Union (mIoU) exceeds 0.75. A reward score of 1.0 is assigned only when both conditions are satisfied; otherwise, the reward is 0. Note that we leave out the common format reward as after the SFT stage, the model has learned to follow the format requirement.
5 Experiments
Training Paradigm Open-ended Closed-ended Single-choice Multi-choice
Dataset SFT RFT A-score† V-score A-score V-score A-score V-score A-score V-score A-score.Avg
Zero-shot - 85.13 12.31 30.39 20.23 49.00 27.81 12.02 7.59 44.14
Fine-tuning GEMeX-Full - 97.31 60.51 82.32 69.32 83.08 67.38 74.51 55.98 84.31
GEMeX-200K - 92.54 56.59 79.93 66.04 78.92 61.39 67.93 49.71 79.83
GEMeX-ThinkVG-200K 94.71 56.52 87.11 66.23 80.08 61.76 70.20 51.13 83.03
GEMeX-ThinkVG-200K 96.44 59.63 90.97 68.39 83.31 64.52 73.18 54.49 85.98
[…]
5.1 Dataset
We conduct experiments on the GEMeX dataset and randomly sample a portion of its training set to generate ThinkVG data. We denote the version with the complete training set of 1.59M VQA triples as GEMeX-Full. The 200K subset that includes ThinkVG data is denoted as GEMeX-ThinkVG-200K, while the subset without ThinkVG is referred to as GEMeX-200K. For evaluation, we use the original GEMeX test set, which consists of 300 images and 3,960 questions (1,144 open-ended questions, 543 closed-ended questions, 1,300 single-choice questions, and 973 multiple-choice questions).
5.2 Training Details
We mainly explore the Qwen2.5-VL-7B-Instruct [^2] https://huggingface.co/Qwen/Qwen2.5-VL-7B-Instruct for experimental verification. During the SFT stage, we fine-tune both the visual projection layers and the LLM components by jointly predicting ThinkVG paths and task-related results. Particularly, the model is trained for 2 epochs on 8 NVIDIA H100 GPUs with a batch size of 256. The network is warmed up in the first 0.05 epochs with a linear learning rate from 3e-7 to 1e-4, which further decays by cosine schedule. The optimizer is AdamW. To accelerate training, we employ the DeepSpeed ZeRO-2 and the bf16 (Brain Floating Point) data format. During the RFT stage, the model is trained for 1 epoch on 8 NVIDIA H100 GPUs with a batch size of 128. The optimizer is AdamW with a learning rate of $62e-$ 6. Similar to SFT, we also fine-tune both the visual projector and LLM components. Regarding the hyper-parameters in Equation 2, we generate $8$ outputs (i.e., $G$) for each input and set $\beta$ and $\epsilon$ to $31e-$ 3 and $0.2$, respectively. The sample generator and LLM used for semantics reward are deployed by vLLM [14] framework.
5.3 Main Results
We report the main results in Table 4. It can be seen that (1) Compared to zero-shot learning, fine-tuning can significantly improve the model’s performance; (2) When SFT is conducted with only 200K vanilla data (i.e., GEMeX-200K), the model lags behind the one trained on the full dataset by a large margin, with an average A-score that is around 4.5% lower; (3) When ThinkVG data (GEMeX-ThinkVG-200K) is further used for SFT, the model’s performance improves substantially, even surpassing that of the model trained on the full dataset in certain tasks, e.g., an improvement of approximately 5% on closed-ended tasks. Nevertheless, it can be observed that when only SFT is used, the improvement in visual grounding is quite limited. This is often due to inaccurate location reasoning in the thinking path after learning ThinkVG with SFT, which leads to error propagation in the final outputs; (4) To address this issue, further applying RFT with semantics reward and grounding reward yields noticeable enhancement, a 1.5%–3% improvement in both answer accuracy and localization performance. It is worth noting that, compared to using the full training data, the model further fine-tuned with RFT achieves better results in answer generation (i.e., higher average A-score) and comparable performance in localization. These findings demonstrate the effectiveness of our proposed ThinkVG dataset and RFT rewards.
Task SFT Data Original A-score Per.1 Per.2 Per.3 Std.†
Closed. 200K 79.93 76.98 83.06 75.69 3.214
ThinkVG-200K 87.11 88.39 86.56 87.66 0.752
Single. 200K 78.92 79.00 77.69 77.31 0.723
ThinkVG-200K 80.08 80.15 80.00 79.69 0.192
[…]
5.4 Robustness to Input Variations
We conduct two common perturbations to generate input variations: (1) we modify closed-ended questions while preserving their semantic integrity—for example, e.g., changing “Is the heart size abnormal in this CXR?” by randomly replacing the interrogative word “is” with “isn’t”, or replacing “abnormal” with “normal”, and adjusting the corresponding answer accordingly. During the test phase, each question has a 50% chance of being modified by one of the two modes mentioned above. If a question does not meet the criteria for either mode, no modification will be applied; (2) we randomly change the order of the options in single-choice questions. For each type of perturbation, we conduct three rounds of test and show models’ performance in Table 5. To more clearly compare the effectiveness of ThinkVG, we compare the performance of the models only after SFT. The results show that after fine-tuning the model with ThinkVG, its robustness against interference is significantly improved (i.e., lower standard deviation), highlighting another advantage.
5.5 Qualitative Evaluation
As shown in Figure 4 and 5, we compare the performance of models on several challenging examples, trained with and without ThinkVG on GEMeX-200K. When using ThinkVG, we present the model outputs after RFT stage. It can be seen that when using ThinkVG with proper tuning, not only are the answers more accurate, but the model also carefully thinks with analyzing corresponding regions in the image when addressing questions, leading to stronger overall interpretability.
Figure 4: Challenging examples from GEMeX answered by models trained with and without ThinkVG. (✓) or (✗) in outputs highlight correct or incorrect reasons or answers. The colored words indicate the thinking with visual grounding process.
Figure 5: Challenging examples from GEMeX answered by models trained with and without ThinkVG. (✓) or (✗) in outputs highlight correct or incorrect reasons or answers. The colored words indicate the thinking with visual grounding process.
6 Conclusion
In this work, we address the critical issue of interpretability in Medical Visual Question Answering by introducing a thinking with visual grounding dataset (ThinkVG). It explicitly links thinking steps to corresponding regions in the medical image, offering fine-grained visual explanations that enhance transparency and trust in model-generated answers. To further guide the model in generating accurate thinking paths, we proposed a new reward mechanism for reinforcement learning that aligns the model’s outputs with relevant visual and textual evidence. Despite using a subset of training data, the trained model is able to achieve impressive performance. We believe this work marks a step toward making Medical VQA systems more explainable and clinically usable, paving the way for safer and more reliable AI-assisted diagnosis.